# A Survey on Self-Evolution of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have significantly advanced in various fields and intelligent agent applications. However, current LLMs that learn from human or external model supervision are costly and may face performance ceilings as task complexity and diversity increase. To address this issue, self-evolution approaches that enable LLM to autonomously acquire, refine, and learn from experiences generated by the model itself are rapidly growing. This new training paradigm inspired by the human experiential learning process offers the potential to scale LLMs towards superintelligence. In this work, we present a comprehensive survey of self-evolution approaches in LLMs. We first propose a conceptual framework for self-evolution and outline the evolving process as iterative cycles composed of four phases: experience acquisition, experience refinement, updating, and evaluation. Second, we categorize the evolution objectives of LLMs and LLM-based agents; then, we summarize the literature and provide taxonomy and insights for each module. Lastly, we pinpoint existing challenges and propose future directions to improve self-evolution frameworks, equipping researchers with critical insights to fast-track the development of self-evolving LLMs. Our corresponding GitHub repository is available at https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/Awesome-Self-Evolution-of-LLM.
A Survey on Self-Evolution of Large Language Models
Zhengwei Tao12thanks: Work done while interning at Alibaba Group., Ting-En Lin2, Xiancai Chen1, Hangyu Li2, Yuchuan Wu2, Yongbin Li2thanks: Corresponding authors.,  Zhi Jin1†, Fei Huang2, Dacheng Tao3, Jingren Zhou2 1 Key Lab of HCST (PKU), MOE; School of Computer Science, Peking University 2Alibaba Group  3Nanyang Technological University {tttzw, xiancaich}@stu.pku.edu.cn,  zhijin@pku.edu.cn {ting-en.lte, shengxiu.wyc, shuide.lyb, jingren.zhou}@alibaba-inc.com dacheng.tao@ntu.edu.sg
@@ A Survey on Self-Evolution of Large Language Models
1 Introduction
With the rapid development of artificial intelligence, large language models (LLMs) like GPT-3.5 Ouyang et al. (2022), GPT-4 Achiam et al. (2023), Gemini Team et al. (2023), LLaMA Touvron et al. (2023a, b), and Qwen Bai et al. (2023) mark a significant shift in language understanding and generation. These models undergo three stages of development as shown in Figure 1: pre-training on large and diverse corpora to gain a general understanding of language and world knowledge Devlin et al. (2018); Brown et al. (2020), followed by supervised fine-tuning to elicit the abilities of downstream tasks Raffel et al. (2020); Chung et al. (2022). Finally, the human preference alignment training enables the LLMs to respond as human behaviors Ouyang et al. (2022). Such successive training paradigms achieve significant breakthroughs, enabling LLMs to perform a wide range of tasks with remarkable zero-shot and in-context capabilities, such as question answering Tan et al. (2023), mathematical reasoning Collins et al. (2023), code generation Liu et al. (2024b), and task-solving that require interaction with environments Liu et al. (2023b).
Despite these advancements, humans anticipate that the emerging generation of LLMs can be tasked with assignments of greater complexity, such as scientific discovery Miret and Krishnan (2024) and future events forecasting Schoenegger et al. (2024). However, current LLMs encounter challenges in these sophisticated tasks due to the inherent difficulties in modeling, annotation, and the evaluation associated with existing training paradigms Burns et al. (2023). Furthermore, the recently developed Llama-3 model has been trained on an extensive corpus comprising 15 trillion tokens [^1] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct. It’s a monumental volume of data, suggesting that significantly scaling model performance by adding more real-world data could pose a limitation. This has attracted interest in self-evolving mechanisms for LLMs, akin to the natural evolution of human intelligence and illustrated by AI developments in gaming, such as the transition from AlphaGo Silver et al. (2016) to AlphaZero Silver et al. (2017). AlphaZero’s self-play method, requiring no labeled data, showcases a path forward for LLMs to surpass current limitations and achieve superhuman performance without intensive human supervision.
Drawing inspiration from the paradigm above, research on the self-evolution of LLMs has rapidly increased at different stages of model development, such as self-instruct Wang et al. (2023b), self-play Tu et al. (2024), self-improving Huang et al. (2022), and self-training Gulcehre et al. (2023). Notably, DeepMind’s AMIE system Tu et al. (2024) outperforms primary care physicians in diagnostic accuracy, and Microsoft’s WizardLM-2  [^2] https://wizardlm.github.io/WizardLM2/ exceeds the performance of the initial version of GPT-4. Both models are developed using self-evolutionary frameworks with autonomous learning capabilities and represent a potential LLM training paradigm shift. However, the relationships between these methods remain unclear, lacking systematic organization and analysis.
@@ 2.1 Background
Self-Evolution in Artificial Intelligence. Artificial Intelligence represents an advanced form of intelligent agent, equipped with cognitive faculties and behaviors mirroring those of humans. The aspiration of AI developers lies in enabling AI to harness self-evolutionary capabilities, paralleling the experiential learning processes characteristic of human development. The concept of self-evolution in AI emerges from the broader fields of machine learning and evolutionary algorithms Bäck and Schwefel (1993). Initially influenced by the principles of natural evolution, such as selection, mutation, and reproduction, researchers have developed algorithms that simulate these processes to optimize solutions to complex problems. The landmark paper by Holland (1992), which introduced the genetic algorithm, marks a foundational moment in the history of AI’s capability for self-evolution. Subsequent developments in neural networks and deep learning have furthered this capability, allowing AI systems to modify their own architectures and improve performance without human intervention Liu et al. (2021).
Can Artificial Entities Evolve Themselves? Philosophically, the question of whether artificial entities can self-evolve touches on issues of autonomy, consciousness, and agency. While some philosophers argue that true self-evolution in AI would require some form of consciousness or self-awareness, others maintain that mechanical self-improvement through algorithms does not constitute genuine evolution Chalmers (1997). This debate often references the works of thinkers like Dennett (1993), who explore the cognitive processes under human consciousness and contrast them with artificial systems. Ultimately, the philosophical inquiry into AI’s capacity for self-evolution remains deeply intertwined with interpretations of what it means to ’evolve’ and whether such processes can purely be algorithmic or must involve emergent consciousness Searle (1986).
@@ 2.2 Conceptual Framework
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, [Self-Evolution of LLMs, root, parent anchor=south [Evaluation (§7), node_lv1 [Qualitative, node_lv2, [Yang et al. (2023b) , LLM Explanation Zheng et al. (2024a) , ChatEval Chan et al. (2023) , node_lv3], ] [Quantitative, node_lv2, [ LLM-as-a-Judge Zheng et al. (2024a); Dubois et al. (2024) , Reward Score Ouyang et al. (2022) , node_lv3], ] ] [Updating (§6), node_lv1, [In-Context, node_lv2 [Working Memory: Reflexion Shinn et al. (2023), IML Wang et al. (2024a), EvolutionaryAgent Li et al. (2024c) , Agent-Pro Zhang et al. (2024d), ProAgent Zhang et al. (2024a) , node_lv3], [External Memory: MoT Li and Qiu (2023), MemoryBank Zhong et al. (2024b), TiM Liu et al. (2023a), IML Wang et al. (2024a), TRAN Yang et al. (2023b), MemGPT Packer et al. (2023) UA2 Yang et al. (2024d) , ICE Qian et al. (2024), AesopAgent Wang et al. (2024d) , node_lv3], ] [In-Weight, node_lv2 [Architecture: LoRA Hu et al. (2021), ConPET Song et al. (2023), Model Soups Wortsman et al. (2022), DAIR Yu et al. (2023a), UltraFuser Ding et al. (2024), EvoLLM Akiba et al. (2024) , node_lv3], [Regularization: InstuctGPT Ouyang et al. (2022), FuseLLM Wan et al. (2024) , Elastic Reset Noukhovitch et al. (2024), WARM Ramé et al. (2024), AMA Lin et al. (2024) , node_lv3], [Replay: RFT Yuan et al. (2023), ReST Gulcehre et al. (2023); Aksitov et al. (2023) , AMIE Tu et al. (2024), SOTOPIA-$\pi$ Wang et al. (2024e), LLM2LLM Lee et al. (2024), LTC Wang et al. (2023a), A3T Yang et al. (2024c), SSR Huang et al. (2024a), SDFT Yang et al. (2024b) , node_lv3], ] ] [Experience Refinement (§5), node_lv1, [Correcting, node_lv2, [Critique-Free: STaR Zelikman et al. (2022), Self-Debugging Chen et al. (2023c), IterRefinement Chen et al. (2023b), Clinical SV Gero et al. (2023) , node_lv3], [Critique-Based: Self-Refine Madaan et al. (2023), CAI Bai et al. (2022), RCI Kim et al. (2023) , SELF Lu et al. (2023), CRITIC Gou et al. (2023), SelfEvolve Jiang et al. (2023), ISR-LLM Zhou et al. (2023b), Reflexion Shinn et al. (2024) , node_lv3], ] [Filtering, node_lv2, [Metric-Free: Self-Consistency Wang et al. (2022) , LMSI Huang et al. (2022), Self-Verification Weng et al. (2023) , CodeT Chen et al. (2022) , node_lv3] [Metric-Based: ReSTEM Singh et al. (2023), AutoAct Qiao et al. (2024), Self-Talk Ulmer et al. (2024), Self-Instruct Wang et al. (2023b) , node_lv3], ] ] [Experience Acquisition (§4), node_lv1 , [Feedback (§4.3), node_lv1_5 [Environment, node_lv2, [ SelfEvolve Jiang et al. (2023), Self-Debugging Chen et al. (2023c), Reflexion Shinn et al. (2023), CRITIC Gou et al. (2023), RoboCat Bousmalis et al. (2023), SinViG Xu et al. (2024b), SOTOPIA-$\pi$ Wang et al. (2024e) , node_lv3_5], ] [Model, node_lv2, [ Self-Reward Yuan et al. (2024), LSX Stammer et al. (2023), DLMA Liu et al. (2024a), SIRLC Pang et al. (2023), Self-Alignment Zhang et al. (2024e), CAI Bai et al. (2022), Self-Refine Madaan et al. (2023) , node_lv3_5], ] ] [Solution (§4.2), node_lv1_5 [Negative, node_lv2, [Perturbative: RLCD Yang et al. (2023a), DLMA Liu et al. (2024a), Ditto Lu et al. (2024a) , node_lv3_5], [Contrastive: Self-Reward Yuan et al. (2024), SPIN Chen et al. (2024), GRATH Chen and Li (2024), Self-Contrast Zhang et al. (2024c), ETO Song et al. (2024), A3Yang et al. (2024c), STE Wang et al. (2024b), COTERRORSET Tong et al. (2024) , node_lv3_5], ] [Positive, node_lv2, [Grounded: Self-Align Sun et al. (2024), SALMON Sun et al. (2023), MemoryBank Zhong et al. (2024b), TiM Liu et al. (2023a), MoT Li and Qiu (2023), IML Wang et al. (2024a), TRAN Yang et al. (2023b), MemGPT Packer et al. (2023) , node_lv3_5 ], [Self-Play: Debates Taubenfeld et al. (2024), Self-Talk Ulmer et al. (2024), Ditto Lu et al. (2024a), SOLID Askari et al. (2024), SOTOPIA-$\pi$ Wang et al. (2024e) , node_lv3_5], [Interactive: SelfEvolve Jiang et al. (2023), LDB Zhong et al. (2024a), ETO Song et al. (2024), A3T Yang et al. (2024c), AutoAct Qiao et al. (2024), KnowAgent Zhu et al. (2024) , node_lv3_5], [Rationale-Based: LMSI Huang et al. (2022), STaR Zelikman et al. (2022), A3T Yang et al. (2024c) , node_lv3_5], ] ] [Task (§4.1), node_lv1_5 [Selective, node_lv2, [ DIVERSE-EVOL Wu et al. (2023), SOFT Wang et al. (2024c), Selective Reflection-Tuning Li et al. (2024b) V-STaR Hosseini et al. (2024) , node_lv3_5], ] [Knowledge-Free, node_lv2, [ Self-Instruct Wang et al. (2023b); Honovich et al. (2022); Roziere et al. (2023), Ada-Instruct Cui and Wang (2023), Evol-Instruct Xu et al. (2024a), MetaMath Yu et al. (2023b), PromptBreeder Fernando et al. (2023), Backtranslation Li et al. (2023b), Kun Zheng et al. (2024b) , node_lv3_5], ] [Knowledge-Based, node_lv2, [Unstructured: UltraChat Ding et al. (2023), SciGLM Zhang et al. (2024b), EvIT Tao et al. (2024a), MEEL Tao et al. (2024b) , node_lv3_5], [Structured: Self-Align Sun et al. (2024), Ditto Lu et al. (2024a), SOLID Askari et al. (2024) , node_lv3_5], ] ] ] ]
@@ 3 Evolution Objectives
Evolution objectives in self-evolving LLMs serve as predefined goals that autonomously guide their development and refinement. Much like humans set personal objectives based on needs and desires, these objectives are crucial as they determine how the model iteratively self-updates. They enable the LLM to autonomously learn from new data, optimize algorithms, and adapt to changing environments, effectively "feeling" its needs from feedback or self-assessment and setting its own goals to enhance functionality without human intervention.
@@
where ${\mathcal{E}}^{t}$ is the evolution objective, composed by evolving abilities ${\mathcal{A}}^{t}$ and evolution directions ${\mathcal{D}}^{t}$. Take "reasoning accuracy improving" as an example, "reasoning" is the evolving ability and "accuracy improving" is the evolution direction.

### v2

@@ Abstract
Large language models (LLMs) have significantly advanced in various fields and intelligent agent applications. However, current LLMs that learn from human or external model supervision are costly and may face performance ceilings as task complexity and diversity increase. To address this issue, self-evolution approaches that enable LLM to autonomously acquire, refine, and learn from experiences generated by the model itself are rapidly growing. This new training paradigm inspired by the human experiential learning process offers the potential to scale LLMs towards superintelligence. In this work, we present a comprehensive survey of self-evolution approaches in LLMs. We first propose a conceptual framework for self-evolution and outline the evolving process as iterative cycles composed of four phases: experience acquisition, experience refinement, updating, and evaluation. Second, we categorize the evolution objectives of LLMs and LLM-based agents; then, we summarize the literature and provide taxonomy and insights for each module. Lastly, we pinpoint existing challenges and propose future directions to improve self-evolution frameworks, equipping researchers with critical insights to fast-track the development of self-evolving LLMs. Our corresponding GitHub repository is available at https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/Awesome-Self-Evolution-of-LLM. 00footnotetext: Work done while Zhengwei Tao was interning at Alibaba Group.
keywords:
Large Language Model, Self-Evolution, Self-Improvement, Self-Training, Autonomous Agents
1 Introduction
With the rapid development of artificial intelligence, large language models (LLMs) like GPT-3.5 [83], GPT-4 [1], Gemini [113], LLaMA [115, 116], and Qwen [8] mark a significant shift in language understanding and generation. These models undergo three stages of development as shown in Figure 1: pre-training on large and diverse corpora to gain a general understanding of language and world knowledge [30, 14], followed by supervised fine-tuning to elicit the abilities of downstream tasks [91, 25]. Finally, the human preference alignment training enables the LLMs to respond as human behaviors [83]. Such successive training paradigms achieve significant breakthroughs, enabling LLMs to perform a wide range of tasks with remarkable zero-shot and in-context capabilities, such as question answering [109], mathematical reasoning [26], code generation [71], and task-solving that require interaction with environments [73].
@@ 1 Introduction
Despite these advancements, humans anticipate that the emerging generation of LLMs can be tasked with assignments of greater complexity, such as scientific discovery [81] and future events forecasting [95]. However, current LLMs encounter challenges in these sophisticated tasks due to the inherent difficulties in modeling, annotation, and the evaluation associated with existing training paradigms [15]. Furthermore, the recently developed Llama-3 model has been trained on an extensive corpus comprising 15 trillion tokens [^1] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct. It’s a monumental volume of data, suggesting that significantly scaling model performance by adding more real-world data could pose a limitation.
This has attracted interest in self-evolving mechanisms for LLMs, akin to the natural evolution of human intelligence and illustrated by AI developments in gaming, such as the transition from AlphaGo [101] to AlphaZero [102]. AlphaZero’s self-play method, requiring no labeled data, showcases a path forward for LLMs to surpass current limitations and achieve superhuman performance without intensive human supervision.
Drawing inspiration from the paradigm above, research on the self-evolution of LLMs has rapidly increased at different stages of model development, such as self-instruct [127], self-play [117], self-improving [52], and self-training [42]. Notably, DeepMind’s AMIE system [117] outperforms primary care physicians in diagnostic accuracy, and Microsoft’s WizardLM-2  [^2] https://wizardlm.github.io/WizardLM2/ exceeds the performance of the initial version of GPT-4. Both models are developed using self-evolutionary frameworks with autonomous learning capabilities and represent a potential LLM training paradigm shift. However, the relationships between these methods remain unclear, lacking systematic organization and analysis.
@@ 2.1 Background
Self-Evolution in Artificial Intelligence. Artificial Intelligence represents an advanced form of intelligent agent, equipped with cognitive faculties and behaviors mirroring those of humans. The aspiration of AI developers lies in enabling AI to harness self-evolutionary capabilities, paralleling the experiential learning processes characteristic of human development. The concept of self-evolution in AI emerges from the broader fields of machine learning and evolutionary algorithms [7]. Initially influenced by the principles of natural evolution, such as selection, mutation, and reproduction, researchers have developed algorithms that simulate these processes to optimize solutions to complex problems. The landmark paper by Holland [47], which introduced the genetic algorithm, marks a foundational moment in the history of AI’s capability for self-evolution. Subsequent developments in neural networks and deep learning have furthered this capability, allowing AI systems to modify their own architectures and improve performance without human intervention [74].
Can Artificial Entities Evolve Themselves? Philosophically, the question of whether artificial entities can self-evolve touches on issues of autonomy, consciousness, and agency. While some philosophers argue that true self-evolution in AI would require some form of consciousness or self-awareness, others maintain that mechanical self-improvement through algorithms does not constitute genuine evolution [16]. This debate often references the works of thinkers like Dennett [28], who explore the cognitive processes under human consciousness and contrast them with artificial systems. Ultimately, the philosophical inquiry into AI’s capacity for self-evolution remains deeply intertwined with interpretations of what it means to ’evolve’ and whether such processes can purely be algorithmic or must involve emergent consciousness [97].
@@ 2.2 Conceptual Framework
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, [Self-Evolution of LLMs, root, parent anchor=south [Evaluation (§7), node_lv1 [Qualitative, node_lv2, [Yang et al [140] , LLM Explanation [157] , ChatEval [17] , node_lv3], ] [Quantitative, node_lv2, [ LLM-as-a-Judge [157, 34] , Reward Score [83] , node_lv3], ] ] [Updating (§6), node_lv1, [In-Context, node_lv2 [Working Memory: Reflexion [98], IML [121], EvolutionaryAgent [65] , Agent-Pro [154], ProAgent [151] , node_lv3], [External Memory: MoT [66], MemoryBank [161], TiM [72], IML [121], TRAN [140], MemGPT [84] UA2 [142] , ICE [89], AesopAgent [123] , node_lv3], ] [In-Weight, node_lv2 [Architecture: LoRA [50], ConPET [104], Model Soups [130], DAIR [147], UltraFuser [33], EvoLLM [3] , node_lv3], [Regularization: InstuctGPT [83], FuseLLM [119] , Elastic Reset [82], WARM [92], AMA [69] , node_lv3], [Replay: ReST [42, 4] , AMIE [117], SOTOPIA-$\pi$ [125], LLM2LLM [60], LTC [124], A3T [141], SSR [53], SDFT [143] , node_lv3], ] ] [Experience Refinement (§5), node_lv1, [Correcting, node_lv2, [Critique-Free: STaR [149], Self-Debugging [22], IterRefinement [20], Clinical SV [40] , node_lv3], [Critique-Based: Self-Refine [80], CAI [9], RCI [57] , SELF [76], CRITIC [41], SelfEvolve [56], ISR-LLM [164], Reflexion [98] , node_lv3], ] [Filtering, node_lv2, [Metric-Free: Self-Consistency [126] , LMSI [52], Self-Verification [129] , CodeT [18] , node_lv3] [Metric-Based: ReSTEM [103], AutoAct [90], Self-Talk [118], Self-Instruct [127] , node_lv3], ] ] [Experience Acquisition (§4), node_lv1 , [Feedback (§4.3), node_lv1_5 [Environment, node_lv2, [ SelfEvolve [56], Self-Debugging [22], Reflexion [98], CRITIC [41], RoboCat [13], SinViG [135], SOTOPIA-$\pi$ [125] , node_lv3_5], ] [Model, node_lv2, [ Self-Reward [148], LSX [106], DLMA [70], SIRLC [85], Self-Alignment [155], CAI [9], Self-Refine [80] , node_lv3_5], ] ] [Solution (§4.2), node_lv1_5 [Negative, node_lv2, [Perturbative: RLCD [138], DLMA [70], Ditto [77] , node_lv3_5], [Contrastive: Self-Reward [148], SPIN [24], GRATH [21], Self-Contrast [153], ETO [105], A3[141], STE [120], COTERRORSET [114] , node_lv3_5], ] [Positive, node_lv2, [Grounded: Self-Align [108], SALMON [107], MemoryBank [161], TiM [72], MoT [66], IML [121], TRAN [140], MemGPT [84] , node_lv3_5 ], [Self-Play: Debates [112], Self-Talk [118], Ditto [77], SOLID [6], SOTOPIA-$\pi$ [125] , node_lv3_5], [Interactive: SelfEvolve [56], LDB [159], ETO [105], A3T [141], AutoAct [90], KnowAgent [165] , node_lv3_5], [Rationale-Based: LMSI [52], STaR [149], A3T [141] , node_lv3_5], ] ] [Task (§4.1), node_lv1_5 [Selective, node_lv2, [ DIVERSE-EVOL [131], SOFT [122], Selective Reflection-Tuning [64], V-STaR [49] , node_lv3_5], ] [Knowledge-Free, node_lv2, [ Self-Instruct [127, 48, 93], Ada-Instruct [27], Evol-Instruct [134], MetaMath [146], PromptBreeder [36], Backtranslation [67], Kun [158] , node_lv3_5], ] [Knowledge-Based, node_lv2, [Unstructured: UltraChat [32], SciGLM [152], EvIT [110], MEEL [111] , node_lv3_5], [Structured: Self-Align [108], Ditto [77], SOLID [6] , node_lv3_5], ] ] ] ]
@@ 3 Evolution Objectives
Evolution objectives in self-evolving LLMs serve as predefined goals that autonomously guide their development and refinement. Much like humans set personal objectives based on needs and desires, these objectives are crucial as they determine how the model iteratively self-updates. They enable the LLM to autonomously learn from new data, optimize algorithms, and adapt to changing environments, effectively ”feeling” its needs from feedback or self-assessment and setting its own goals to enhance functionality without human intervention.
@@
where ${\mathcal{E}}^{t}$ is the evolution objective, composed by evolving abilities ${\mathcal{A}}^{t}$ and evolution directions ${\mathcal{D}}^{t}$. Take ”reasoning accuracy improving” as an example, ”reasoning” is the evolving ability and ”accuracy improving” is the evolution direction.
@@ 3.1 Evolving Abilities
Task $f^{{\mathcal{T}}}$ Solution $f^{{\mathcal{Y}}}$ Feedback $f^{{\mathcal{F}}}$
Large Language Models
Self-Align [108] Knowledge-Based Pos-G - Filtering In-W IF
SciGLM [152] Knowledge-Based - - - In-W Other
EvIT [110] Knowledge-Based - - - In-W Reasoning
MEEL [111] Knowledge-Based - - - In-W Reasoning
UltraChat [32] Knowledge-Based - - - In-W Role-Play
SOLID [6] Knowledge-Based Pos-S - Filtering In-W Role-Play
Ditto [77] Knowledge-Based Pos-S, Neg-P - - In-W Role-Play
MetaMath [146] Knowledge-Free Pos-R - - In-W Math
Self-Rewarding [148] Knowledge-Free - Model - In-W IF,Reasoning,Role-Play
Kun [158] Knowledge-Free - - Filtering In-W IF,Reasoning
PromptBreeder [36] Knowledge-Free - - - In-C Math, Reasoning
Ada-Instruct [27] Knowledge-Free - - - In-W Math, Reasoning, Code
Backtranslation [67] Knowledge-Free - - - In-W IF
DiverseEvol [131] Selective - - - In-W Code
Grath [21] Selective Neg-C Model - In-W Reasoning
RESTem [103] Selective - Model Filtering In-W Math, Code
SOFT [122] Selective - - - In-W IF
LSX [106] - Pos-R Model Correcting In-W Other
LMSI [52] - Pos-R - Filtering In-W Math
TRAN [140] - Pos-G - - In-C Reasoning
MOT [66] - Pos-R, Pos-G - Filtering In-C Math, Reasoning
STaR [149] - Pos-R, Neg-C Model Correct In-W Reasoning
COTERRORSET [114] - Pos-R, Neg-C - - In-W Math, Reasoning
Self-Debugging [22] - Pos-I Env - In-C Code
SelfEvolve [56] - Pos-I Env - In-C Code
Reflexion [98] - Pos-I, Pos-G - - In-C Code, Reasoning
V-STaR [49] - Neg-C Model Filter In-W Math, Code
Self-Contrast [155] - Neg-C Model - In-W Reasoning
SALMON [107] - Neg-C Model - In-W IF,Reasoning,Role-Play
SPIN [24] - Neg-C - - In-W IF,Reasoning,Role-Play
RLCD [138] - Neg-P Model - In-W IF
DLMA [70] - Neg-P Model - In-W IF
SELF [76] - - Model Correct In-W IF, Math
LLM Agents
AutoAct [90] Knowledge-Based Pos-I Env Filtering In-W Planning, Tool
KnowAgent [165] Knowledge-Based Pos-I, Pos-G Env Filtering In-W Embodied, Planning, Tool
RoboCat [13] Knowledge-Free Pos-I Env - In-W Embodied
STE [120] Knowledge-Free Pos-I, Neg-C Env Correct In-W Tool
IML [121] - Pos-R, Pos-G - - In-C Reasoning
SinViG Xu et al [135] - Pos-I Env Filtering In-W Embodied
ETO [105] - Pos-I, Neg-C Env Correct In-W Tool
A3[141] - Pos-I, Neg-C Env Correct In-W Tool
Debates [112] - Pos-S - - In-W Communication
SOTOPIA-$\pi$ [125] - Pos-S,Pos-G Env - In-W Communication
Self-Talk [118] - Pos-S, Pos-G Model Filtering In-W Communication
MemGPT [84] - Pos-G Env Filtering In-C Communication
MemoryBank [161] - Pos-G Env Filtering In-C Communication
ProAgent [151] - Pos-G Env - In-C Embodied
Agent-Pro [154] - Pos-G Env - In-C Planning
AesopAgent [123] - Pos-G Env - In-C Planning
ICE [89] - Pos-G Env - In-C Planning
TiM [72] - Pos-G - - In-C Communication
Werewolf [136] - Pos-G - - In-C Planning
Table 1: Overview of self-evolution methods, detailing approaches across evolutionary stages. Key: Pos (Positive), Neg (Negative), R (Rationale-based), I (Interactive), S (Self-play), G (Grounded), C (Contrastive), P (Perturbative), Env (Environment), In-W (In-Weight), In-C (In-Context), IF (Instruction-Following). We use colors to highlight evolution directions: Adaptation to Feedback in green, Expansion of Knowledge Base in blue, and Safety, Ethics, and Bias Reduction in brown. Performance enhancement remains in black.